# IFT-UAM/CSIC-25-108

### v1

@@ I Variation in chi-squared weights and model-selection criteria
$\chi\mathrm{SU(3)}][a^{2}][-]$ 0.84 0.0405 0.1446(7)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40]$ 0.77 0.0226 0.1446(9)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.46]$ 0.68 0.0117 0.1443(11)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.77 0.0188 0.1443(7)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.53 0.0059 0.1441(13)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<350\,\mathrm{MeV}]$ 0.60 0.0037 0.1446(8)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}L>4.1]$ 0.91 0.0223 0.1447(8)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.85 0.0404 0.1445(7)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.77 0.0226 0.1446(9)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.69 0.0118 0.1443(12)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\,\mathrm{MeV}]$ 0.77 0.0188 0.1442(7)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.55 0.0059 0.1441(13)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<350\,\mathrm{MeV}]$ 0.61 0.0037 0.1446(8)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.90 0.0223 0.1447(9)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][-]$ 0.83 0.0344 0.1440(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.73 0.0168 0.1440(14)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.60 0.0081 0.1443(20)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.70 0.0128 0.1445(13)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.88 0.0160 0.1440(15)
$\chi\mathrm{SU(2)}][a^{2}][-]$ 0.87 0.0382 0.1441(9)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40]$ 0.79 0.0205 0.1442(11)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.46]$ 0.67 0.0102 0.1440(13)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.74 0.0138 0.1439(10)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.47 0.0035 0.1436(16)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<350\,\mathrm{MeV}]$ 0.91 0.0054 0.1435(13)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}L>4.1]$ 0.84 0.0153 0.1446(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.86 0.0382 0.1441(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.79 0.0205 0.1441(11)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.67 0.0096 0.1440(13)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\,\mathrm{MeV}]$ 0.73 0.0137 0.1439(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.46 0.0033 0.1436(16)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<350\,\mathrm{MeV}]$ 0.90 0.0053 0.1435(13)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.85 0.0153 0.1445(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][-]$ 0.82 0.0257 0.1441(13)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.73 0.0136 0.1442(18)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.59 0.0060 0.1450(26)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.72 0.0087 0.1448(15)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.81 0.0102 0.1438(19)
@@
$\mathrm{Tay2}][a^{2}][-]$ 0.72 0.0223 0.1449(7)
$\mathrm{Tay2}][a^{2}][\beta>3.40]$ 0.63 0.0125 0.1450(9)
$\mathrm{Tay2}][a^{2}][\beta>3.46]$ 0.53 0.0072 0.1447(12)
$\mathrm{Tay2}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.64 0.0112 0.1446(7)
$\mathrm{Tay2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.40 0.0035 0.1445(13)
$\mathrm{Tay2}][a^{2}][m_{\pi}<350\,\mathrm{MeV}]$ 0.43 0.0024 0.1449(8)
$\mathrm{Tay2}][a^{2}][m_{\pi}L>4.1]$ 0.92 0.0243 0.1451(7)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.70 0.0222 0.1449(7)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.63 0.0125 0.1450(10)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.53 0.0072 0.1447(12)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\,\mathrm{MeV}]$ 0.63 0.0112 0.1446(8)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.40 0.0035 0.1445(14)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<350\,\mathrm{MeV}]$ 0.43 0.0023 0.1449(8)
$\mathrm{Tay2}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.91 0.0242 0.1451(8)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][-]$ 0.70 0.0199 0.1443(11)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.60 0.0096 0.1443(16)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.45 0.0043 0.1447(21)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.56 0.0074 0.1448(13)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.92 0.0197 0.1441(16)
$\mathrm{Tay4}][a^{2}][-]$ 0.83 0.0300 0.1441(9)
$\mathrm{Tay4}][a^{2}][\beta>3.40]$ 0.77 0.0166 0.1441(11)
$\mathrm{Tay4}][a^{2}][\beta>3.46]$ 0.64 0.0083 0.1438(13)
$\mathrm{Tay4}][a^{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.71 0.0127 0.1439(9)
$\mathrm{Tay4}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\,\mathrm{MeV}]$ 0.44 0.0031 0.1436(15)
$\mathrm{Tay4}][a^{2}][m_{\pi}<350\,\mathrm{MeV}]$ 0.90 0.0053 0.1435(11)
$\mathrm{Tay4}][a^{2}][m_{\pi}L>4.1]$ 0.87 0.0158 0.1453(14)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][-]$ 0.78 0.0204 0.1441(11)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.70 0.0106 0.1442(15)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.60 0.0056 0.1451(21)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\,\mathrm{MeV}]$ 0.74 0.0094 0.1449(13)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.95 0.0183 0.1442(17)
@@
$\chi\mathrm{SU(3)}][a^{2}][-]$ 0.78 0.0503 0.1448(8)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40]$ 0.69 0.0256 0.1446(11)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.46]$ 0.56 0.0108 0.1448(14)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.75 0.0307 0.1446(9)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.50 0.0069 0.1446(15)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.47 0.0025 0.1446(10)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}L>4.1]$ 0.81 0.0154 0.1449(9)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.77 0.0502 0.1448(8)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.69 0.0255 0.1446(12)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.56 0.0109 0.1448(14)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.75 0.0307 0.1445(9)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.51 0.0069 0.1446(15)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.47 0.0025 0.1446(10)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.82 0.0153 0.1449(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][-]$ 0.77 0.040 0.1443(12)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.65 0.0165 0.1440(19)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.47 0.0058 0.1449(22)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.69 0.0177 0.1445(14)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.79 0.0108 0.1440(19)
$\chi\mathrm{SU(2)}][a^{2}][-]$ 0.79 0.0422 0.1445(9)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40]$ 0.68 0.0184 0.1443(13)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.46]$ 0.53 0.0077 0.1446(15)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.70 0.0152 0.1445(10)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.39 0.0024 0.1446(18)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.96 0.0052 0.1432(12)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}L>4.1]$ 0.74 0.0074 0.1447(12)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}$][-] 0.78 0.0422 0.1445(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40]$ 0.69 0.0183 0.1442(14)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.46]$ 0.54 0.0077 0.1445(15)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.70 0.0151 0.1445(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;{\mathrm{MeV}}]$ 0.39 0.0029 0.1445(18)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}<350\;{\mathrm{MeV}}]$ 0.97 0.0051 0.1432(12)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{\rm\scriptscriptstyle s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.75 0.0087 0.1447(12)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][-]$ 0.73 0.0261 0.1444(14)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.60 0.0089 0.1441(22)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.47 0.0041 0.1460(28)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;{\mathrm{MeV}}]$ 0.60 0.0071 0.1444(17)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.70 0.0055 0.1438(23)

### v2

@@ I Variation in chi-squared weights and model-selection criteria
$\chi\mathrm{SU(3)}][a^{2}][-]$ 0.60 0.0645 0.1437(6)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40]$ 0.52 0.0243 0.1435(8)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.46]$ 0.42 0.0113 0.1432(9)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.51 0.0201 0.1437(6)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.32 0.0042 0.1431(10)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.33 0.0025 0.1438(7)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}L>4.1]$ 0.44 0.0107 0.1439(7)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.58 0.0662 0.1438(6)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.52 0.0237 0.1435(8)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.42 0.0113 0.1433(9)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.50 0.0191 0.1438(6)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.34 0.0042 0.1432(10)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<350\;\rm{MeV}]$ 0.34 0.0025 0.1439(7)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.44 0.0103 0.1440(7)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][-]$ 0.57 0.0430 0.1433(10)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.45 0.0161 0.1432(13)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.33 0.0051 0.1433(17)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.44 0.0103 0.1433(13)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.37 0.0061 0.1437(13)
$\chi\mathrm{SU(2)}][a^{2}][-]$ 0.56 0.0383 0.1438(9)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40]$ 0.47 0.0156 0.1436(11)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.46]$ 0.39 0.0069 0.1434(11)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.45 0.0110 0.1438(10)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.25 0.0021 0.1434(13)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.68 0.0043 0.1423(13)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}L>4.1]$ 0.42 0.0070 0.1442(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.55 0.0365 0.1439(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.45 0.0173 0.1437(11)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.39 0.0065 0.1435(12)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.43 0.0097 0.1439(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.25 0.0021 0.1435(13)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<350\;\rm{MeV}]$ 0.69 0.0044 0.1424(13)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.41 0.0062 0.1443(10)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][-]$ 0.54 0.0275 0.1430(12)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.42 0.0108 0.1429(16)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.29 0.0030 0.1428(20)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.41 0.0065 0.1429(15)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.41 0.0047 0.1430(17)
@@
$\mathrm{Tay2}][a^{2}][-]$ 0.56 0.0572 0.1441(7)
$\mathrm{Tay2}][a^{2}][\beta>3.40]$ 0.48 0.0241 0.1439(9)
$\mathrm{Tay2}][a^{2}][\beta>3.46]$ 0.39 0.0080 0.1436(10)
$\mathrm{Tay2}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.47 0.0170 0.1440(7)
$\mathrm{Tay2}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.30 0.0037 0.1436(11)
$\mathrm{Tay2}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.22 0.0014 0.1441(9)
$\mathrm{Tay2}][a^{2}][m_{\pi}L>4.1]$ 0.51 0.0150 0.1443(8)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.55 0.0567 0.1442(8)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.48 0.0245 0.1440(9)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.39 0.0080 0.1437(10)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.48 0.0163 0.1442(8)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.30 0.0037 0.1437(11)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<350\;\rm{MeV}]$ 0.22 0.0014 0.1442(9)
$\mathrm{Tay2}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.49 0.0142 0.1444(8)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][-]$ 0.55 0.0444 0.1434(11)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.44 0.0158 0.1433(14)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.31 0.0051 0.1434(18)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.42 0.0112 0.1434(14)
$\mathrm{Tay2}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.46 0.0089 0.1436(14)
$\mathrm{Tay4}][a^{2}][-]$ 0.46 0.0223 0.1439(9)
$\mathrm{Tay4}][a^{2}][\beta>3.40]$ 0.37 0.0088 0.1437(11)
$\mathrm{Tay4}][a^{2}][\beta>3.46]$ 0.28 0.0031 0.1435(12)
$\mathrm{Tay4}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.37 0.0066 0.1439(9)
$\mathrm{Tay4}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.20 0.0014 0.1434(12)
$\mathrm{Tay4}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.70 0.0043 0.1423(12)
$\mathrm{Tay4}][a^{2}][m_{\pi}L>4.1]$ 0.43 0.0068 0.1449(12)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][-]$ 0.45 0.0148 0.1434(11)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.34 0.0051 0.1433(14)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.19 0.0018 0.1434(18)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.31 0.0043 0.1434(14)
$\mathrm{Tay4}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.55 0.0081 0.1436(15)
@@
$\chi\mathrm{SU(3)}][a^{2}][-]$ 0.74 0.0601 0.1443(5)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40]$ 0.64 0.0343 0.1441(7)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.46]$ 0.54 0.0098 0.1440(9)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.69 0.0287 0.1441(5)
$\chi\mathrm{SU(3)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.46 0.0074 0.1438(9)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.50 0.0026 0.1444(6)
$\chi\mathrm{SU(3)}][a^{2}][m_{\pi}L>4.1]$ 0.76 0.0123 0.1443(6)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.74 0.0570 0.1443(5)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.65 0.0344 0.1441(7)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.53 0.0099 0.1440(9)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.70 0.0288 0.1441(5)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.44 0.0075 0.1438(9)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<350\;\rm{MeV}]$ 0.49 0.0026 0.1444(6)
$\chi\mathrm{SU(3)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.77 0.0122 0.1443(6)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][-]$ 0.74 0.0495 0.1436(9)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.63 0.0159 0.1434(13)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.44 0.0061 0.1444(18)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.63 0.0147 0.1441(11)
$\chi\mathrm{SU(3)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.75 0.0101 0.1433(14)
$\chi\mathrm{SU(2)}][a^{2}][-]$ 0.74 0.0345 0.1440(9)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40]$ 0.63 0.0168 0.1439(11)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.46]$ 0.48 0.0068 0.1439(12)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<420\;\rm{MeV}]$ 0.63 0.0102 0.1440(10)
$\chi\mathrm{SU(2)}][a^{2}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.32 0.0021 0.1438(13)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}<350\;\rm{MeV}]$ 0.99 0.0046 0.1430(12)
$\chi\mathrm{SU(2)}][a^{2}][m_{\pi}L>4.1]$ 0.68 0.0073 0.1444(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}$][-] 0.73 0.0323 0.1440(9)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40]$ 0.62 0.0169 0.1439(11)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.46]$ 0.48 0.0070 0.1439(12)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<420\;\rm{MeV}]$ 0.64 0.0105 0.1440(10)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][\beta>3.40\;\&\;m_{\pi}<420\;\rm{MeV}]$ 0.32 0.0021 0.1438(13)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}<350\;\rm{MeV}]$ 0.99 0.0046 0.1430(12)
$\chi\mathrm{SU(2)}][a^{2}\alpha_{s}^{\hat{\Gamma}}][m_{\pi}L>4.1]$ 0.68 0.0073 0.1443(11)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][-]$ 0.70 0.0242 0.1435(12)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.40]$ 0.54 0.0086 0.1433(16)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][\beta>3.46]$ 0.38 0.0032 0.1445(23)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}<420\;\rm{MeV}]$ 0.53 0.0056 0.1440(14)
$\chi\mathrm{SU(2)}][a^{2}+a^{2}\phi_{2}][m_{\pi}L>4.1]$ 0.71 0.0064 0.1428(18)